# LiDAR, GNSS and IMU Sensor Alignment through Dynamic Time Warping to Construct 3D City Maps

### v1

@@ Abstract
LiDAR-based 3D mapping suffers from cumulative drift causing global misalignment, particularly in GNSS-constrained environments. To address this, we propose a unified framework that fuses LiDAR, GNSS, and IMU data for high-resolution city-scale mapping. The method performs velocity-based temporal alignment using Dynamic Time Warping and refines GNSS and IMU signals via extended Kalman filtering. Local maps are built using Normal Distributions Transform-based registration and pose graph optimization with loop closure detection, while global consistency is enforced using GNSS-constrained anchors followed by fine registration of overlapping segments. We also introduce a large-scale multimodal dataset captured in Perth, Western Australia to facilitate future research in this direction. Our dataset comprises 144,000 frames acquired with a 128-channel Ouster LiDAR, synchronized RTK-GNSS trajectories, and MEMS-IMU measurements across 21 urban loops. To assess geometric consistency, we evaluated our method using alignment metrics based on road centerlines and intersections to capture both global and local accuracy. Our method reduces the average global alignment error from 3.32 m to 1.24 m, achieving a 61.4% improvement. The constructed high-fidelity map supports a wide range of applications, including smart city planning, geospatial data integration, infrastructure monitoring, and GPS-free navigation. Our method, and dataset together establish a new benchmark for evaluating 3D city mapping in GNSS-constrained environments. The dataset and code will be released publicly.
keywords:
Light Detection & Ranging (LiDAR) , Point Cloud , Global Navigation Satellite System , Inertial Measurement Unit , 3D City Mapping
journal: International Journal of Applied Earth Observation and Geoinformation
[…]
1 Introduction
@@
We evaluate our method using centreline and intersection-based metrics by comparing the reconstructed 3D maps against high-resolution 2D cartographic references  [33, 34]. The proposed method achieves a 61.4% reduction in global alignment error, lowering the RMSE from 3.32m to 1.24m. The modular architecture of the pipeline enables generalization to different sensor configurations and urban layouts, supporting its applicability to broader city-scale mapping scenarios. To promote reproducibility and foster further research, we release our code and dataset.
This work offers the following key contributions:
@@
    LIGMA framework: A unified sensor fusion pipeline that integrates LiDAR, GNSS, and inertial data using a novel velocity-based alignment and multi-stage registration strategy guided by GNSS/IMU priors.
@@
    Urban-scale dataset: A high-resolution dataset covering $4.2\,\text{km}^{2}$ of central Perth, comprising 144,000 LiDAR frames with synchronized and georeferenced inertial and positioning data.
2 Related Work
Precise and globally consistent 3D urban mapping is essential for applications such as infrastructure monitoring, and city-scale planning [35]. Prior work in LiDAR-based mapping has demonstrated reliable geometric accuracy and robustness to lighting conditions, with methods such as GICP [10] and SLIM framework [36, 37] improving global consistency and scalability. Pose graph optimization [38] and semantic-aware registration, including LCDNet [39] and semantic NDT [40], have addressed drift and sparse-feature challenges. However, LiDAR-only approaches like LOAM [41], SliceTransformer [42] and LeGO-LOAM [43] remain limited in feature-scarce or high-speed scenarios due to restricted sensing range and insufficient feature extraction capabilities [44]. This degrades the global map accuracy and continuity, especially when using solid-state LiDARs with limited sensor coverage. Additionally, Dynamic Channel-wise Outlier Removal methods have been adopted to de-noise LiDAR data, enabling the construction of more accurate and consistent point cloud maps [28, 45].
To overcome single-sensor limitations, recent studies increasingly adopt multimodal fusion strategies that combine LiDAR, inertial and positioning sensors for enhanced localization and mapping [30, 17, 29]. IMU integration has been shown to reduce drift to a greater extent compared to LiDAR-only baselines such as NDT or ICP [28]. EKF and factor graph methods are foundational to multi-sensor integration, with the latter providing improved scalability in complex urban environments [30]. Synchronization across heterogeneous sensors has also advanced through velocity-based Dynamic Time Warping, aligning LiDAR-derived velocities with IMU profiles to improve temporal coherence [45, 46]. Despite these advances, many synchronization techniques rely on simplistic interpolation or fixed-rate assumptions, which degrade under real-world jitter and asynchronous sampling [47, 48].
Traditional metrics for 3D map evaluation, including absolute pose error, drift per unit distance, and translation error, remain widely used for assessing global alignment. Translation error quantifies deviations from reference centerlines [49, 33], while drift is typically reported per 100 meters, particularly relevant for loop-free mapping scenarios. Map completion rate has also emerged as a robustness metric, especially in non-revisiting trajectories, with methods like SLIM [10] achieving over 90% accuracy. However, these metrics have critical limitations: they depend on costly ground truth from terrestrial laser scanners, and often fail to accurately capture local geometric structures in dynamic or GNSS-constrained environments [50, 45].
These limitations underscore the need for a unified framework that addresses drift and sensor misalignment challenges in city-scale mapping. Our work introduces such a solution, combining multimodal fusion with a benchmark dataset to real-world constraints.
3 Data Collection
@@
3.1 Survey Objectives and Study Area Definition
@@
3.2 Sensor Specifications
The sensing system comprises a 128-channel Ouster OS1 LiDAR and a RUG-3-IMX-5-Dual GNSS–IMU module [51]. The OS1 is a dual-return high-resolution scanner with a 45° vertical field of view, 0.35° angular resolution at 10 Hz, and a peak output of 2.6 million points per second. Each point includes 3D coordinates, intensity, and a time stamp, which supports detailed reconstruction of urban environments [52, 53]. The GNSS–IMU module integrates dual multi-frequency RTK receivers and a high-performance MEMS inertial unit, delivering sub-meter accuracy and full 6-DoF orientation. It supports GPS, GLONASS, Galileo, and BeiDou constellations. According to the manufacturer specifications, the GNSS operates at 5 Hz and the IMU at 73 Hz [54]. Unlike tightly coupled systems, the LiDAR and GNSS–IMU modules operate asynchronously, with synchronization and logging handled by a central acquisition system as detailed in subsection 3.3.
3.3 Sensor Setup and Hardware Configuration
@@
3.4 Data Modalities, Formats, and Organization
@@
3.5 Quality Assurance and Handling of Imperfections
@@ 4 Methods
4.1 Multimodal Sensor Data Preprocessing
@@
Figure 6: Multimodal data preprocessing pipeline. The top stream illustrates LiDAR data processing: noise filtering, voxel downsampling, and interpolation. The bottom stream shows GNSS/IMU preprocessing: EKF filtering, outlier removal, and smoothing. Outputs include time-synchronized point clouds and trajectories for downstream fusion.
4.1.1 LiDAR Data
Raw LiDAR scans are decoded using the Ouster SDK [55], yielding point cloud frames with 3D coordinates, intensity, reflectivity, and ring ID attributes [52]. To enhance data quality, spurious returns from reflective or transparent surfaces are removed using statistical outlier filtering, which discards points deviating more than one standard deviation ($z=1.0$) from the mean distance within a $k$-nearest neighborhood ($k=50$[44]. For efficiency and spatial consistency, voxel grid downsampling with a 0.05 m resolution is applied, replacing all points in each voxel with their centroid [56, 57]. Frames missing due to occlusion or transmission loss are reconstructed by linear interpolation between adjacent frames [58]. The processed point clouds are stored in .ply format with aligned timestamps for multimodal integration.
4.1.2 GNSS Data
@@
[…]
@@ 4.1.3 IMU Data
4.2 Multimodal Sensor Data Synchronization
Figure 7: Multimodal synchronization pipeline. Velocity profiles from LiDAR, GNSS, and IMU are extracted and aligned. Key temporal features are matched, and Dynamic Time Warping (DTW) resolves temporal offsets. Trimmed outputs yield aligned, time-synchronized sensor data.
@@ 4.2 Multimodal Sensor Data Synchronization
4.2.1 LiDAR Velocity Estimation
To estimate LiDAR velocity, consecutive point cloud frames are registered using ICP algorithm, which yields a 4$\times$4 transformation matrix representing relative motion. The translational displacement $(\Delta x,\Delta y,\Delta z)$ extracted from this matrix is used to compute instantaneous speed as $v_{\text{LiDAR}}(t)=\frac{\sqrt{\Delta x^{2}+\Delta y^{2}+\Delta z^{2}}}{\Delta t}$, where $\Delta t$ is the time interval between frames [63].
4.2.2 IMU Velocity Estimation
@@
[…]
@@
4.2.3 Temporal Alignment with DTW
@@
4.2.4 Temporal Alignment Methodology
@@
$D(i,j)=\left(v_{\text{LiDAR}}(i)-v_{\text{IMU}}(j)\right)^{2}+\min\left\{D(i-1,j),D(i,j-1),D(i-1,j-1)\right\}$ (3)
@@
4.3 High-Precision 3D LiDAR Map Construction
@@
4.3.1 Loop Registration and Pose Graph Optimization
@@
4.3.2 GNSS-Based Loop Calibration
@@
$\mathbf{T}_{\text{loop}}^{\ast}=\arg\min_{\mathbf{T}}\sum_{i=1}^{M}\left\|\mathbf{G}_{\text{loop}}^{(i)}-\mathbf{T}\mathbf{T}_{\text{loop}}^{(i)}\right\|_{\mathbf{W}_{i}}^{2}\vspace{-4mm}$ (5)
@@
4.3.3 Global Merging Using ICP on Overlapping Regions
@@
4.4 Three-Stage Assessment of 3D Map Alignment Accuracy
To assess the spatial accuracy of the proposed LIGMA framework, we implement a structured three-stage protocol (see Fig. 9). First, road centrelines are extracted from the reconstructed 3D LiDAR map and a georeferenced 2D base map from Google Maps, then compared to evaluate geometric alignment. Next, geometric alignment is assessed by computing point-wise correspondences between the LiDAR-derived and reference centrelines, with intersection regions excluded due to ambiguity in lane geometry and resolution. Lastly, intersection zones are analyzed separately using localized deviation metrics to capture alignment accuracy in these geometrically complex areas. These evaluation protocols are described in the following subsections.
Figure 9: Pipeline for quantitative evaluation of 3D map alignment. Road networks extracted from Google Maps are converted to centrelines and registered with LiDAR-derived trajectories, considering left-hand traffic. Alignment accuracy is measured using $\mathcal{L}_{1}$ distance and RMSE, excluding intersections to ensure robustness.
4.4.1 Extraction of Road Centrelines
@@
4.4.2 Centreline-Based Alignment Error Metrics
@@
4.4.3 Intersection-Based Local Geometric Consistency Metrics
@@
$\mathcal{E}_{\text{spatial}}=\frac{1}{K}\sum_{i=1}^{K}\left[\alpha\|p_{i}-m_{i}\|_{2}+\beta\cdot d_{\mathcal{T}}(p_{i},m_{i})\right]\vspace{-4mm}$ (7)
@@
$\mathcal{E}_{\text{env}}=\frac{1}{K}\sum_{i=1}^{K}\max_{x\in\mathcal{B}(p_{i},r)}\min_{y\in\mathcal{B}(m_{i},r)}\|x-y\|_{2}\vspace{-4mm}$ (8)
@@ 5.1 Experimental Setup
5.2 Effectiveness of Multimodal Sensor Data Synchronization
@@
Figure 15: Visualization of 10 selected loops constructed using the traditional LiDAR-only method, showing noticeable geometric distortions.
Figure 16: Visualization of the same 10 loops constructed using our proposed framework, demonstrating improved global alignment and geometric accuracy.
5.3 High-Precision 3D LiDAR Map Construction Results
@@
Table 1: Comparison of road centreline alignment metrics across all test sequences for the LiDAR-only baseline and the proposed framework. Relative improvements indicate percentage reductions in alignment error.
@@
Figure 17: Comparison of road centrelines from the reference map (white), traditional LiDAR-only method (green), and LIGMA (red). The proposed method achieves higher spatial fidelity, closely matching the reference.
Figure 18: Intersection centroid alignment using traditional LiDAR-only. Green points: reference map; red points: LiDAR-only.
Figure 19: Intersection centroid alignment using LIGMA. Green points: reference map; red points: proposed method.
Figure 20: Spatial distribution of 51 urban intersections in Perth CBD. Red markers denote detected intersection centroids, supporting localized geometric analysis across the mapped area.
5.4 Comparative Evaluation of Urban 3D Map Alignment
5.4.1 Visual Comparison of Mapping Results
5.4.2 Loop-Based Centreline Comparison
To highlight quantitative differences, 10 representative loops from a total of 21 are visualized, covering major and geometrically complex road segments in the Perth CBD. Selected for their structural variability and traffic relevance, these loops provide meaningful comparison between methods. Figure 15 shows reconstructions from the traditional LiDAR-only approach. The results exhibit noticeable drift, distortions, and misalignments, especially at intersections and along extended roadways, which lead to irregular loop closures and spatial inconsistencies. In contrast, Figure 16 presents reconstructions using the proposed framework. The resulting trajectories demonstrate significantly improved geometric accuracy and global consistency. Loop shapes are preserved, intersection alignment is enhanced, and drift is substantially reduced, closely matching satellite imagery.
Figure 17 compares centrelines extracted from 3D maps with a 2D reference: white lines denote the reference, green the LiDAR-only method, and red the LIGMA result. The centrelines of our method closely follow the reference, accurately capturing road and intersection geometry. In contrast, the LiDAR-only method exhibits offsets, deformation, and branching artifacts. Table 1 summarizes centreline alignment errors across all 21 sequences, reporting RMSE, minimum, maximum, and standard deviation for both methods. Our proposed framework consistently outperforms the baseline, with percentage reductions across all metrics. On average, LIGMA reduces RMSE from $3.32$ m (LiDAR-only) to $1.24$ m—a $61.4\%$ improvement. Similar gains in minimum, maximum, and standard deviation reflect improved spatial accuracy and robustness across diverse sequences. These enhancements directly support downstream applications such as autonomous navigation and HD map generation.
Table 2: Intersection-level alignment metrics for 15 of 51 intersections(subset shown due to space constraints). CI: Centroidal Offset (m), MOD: Mean Orientation Deviation (°). Abs. = absolute reduction; Pct. = percentage reduction relative to baseline.

### v2

@@ Abstract
LiDAR-based 3D mapping suffers from cumulative drift causing global misalignment, particularly in GNSS-constrained environments. To address this, we propose a unified framework that fuses LiDAR, GNSS, and IMU data for high-resolution city-scale mapping. The method performs velocity-based temporal alignment using Dynamic Time Warping and refines GNSS and IMU signals via extended Kalman filtering. Local maps are built using Normal Distributions Transform-based registration and pose graph optimization with loop closure detection, while global consistency is enforced using GNSS-constrained anchors followed by fine registration of overlapping segments. We also introduce a large-scale multimodal dataset captured in Perth, Western Australia to facilitate future research in this direction. Our dataset comprises 144,000 frames acquired with a 128-channel Ouster LiDAR, synchronized RTK-GNSS trajectories, and MEMS-IMU measurements across 21 urban loops. To assess geometric consistency, we evaluated our method using alignment metrics based on road centerlines and intersections to capture both global and local accuracy. The proposed framework reduces the average global alignment error from 3.32 m to 1.24 m, achieving a 61.4% improvement, and significantly decreases the intersection centroid offset from 13.22 m to 2.01 m, corresponding to an 84.8% enhancement. The constructed high-fidelity map is publicly available through IEEE Dataport and its visualization can be viewed in the provided Demo. This dataset and method together establish a new benchmark for evaluating 3D city mapping in GNSS-constrained environments.
I Introduction
@@
To evaluate geometric consistency, we employ a dual-metric approach using both centreline-based and intersection-based alignment metrics, providing a comprehensive assessment of global and local spatial accuracy. Specifically, our method reduces the global centreline root mean square error (RMSE) from 3.32,m to 1.24,m, representing a 61.4% improvement, and decreases the mean intersection centroid offset from 13.22,m to 2.01,m, achieving an 84.8% enhancement in local geometric fidelity. The complete dataset, consisting of 144,000 LiDAR frames with synchronized RTK-GNSS and MEMS-IMU measurements across 21 urban loops in Perth CBD, is publicly available at IEEE Dataport, with an interactive demonstration accessible at Demo. The modular design of the proposed framework facilitates adaptation to varied sensor configurations and urban topologies, supporting its application in diverse large-scale mapping scenarios and serving as a benchmark for future research in GNSS-challenged urban environments.
Our contributions are summarized below:
@@
    We propose a method for constructing 3D city maps using non-synchrous LiDAR, GNSS and IMU data.
@@
    We propose a unified sensor fusion pipeline, coined LIGMA, that integrates LiDAR, GNSS, and inertial data using a novel velocity-based alignment and multi-stage registration strategy.
  3. 3.
    We present a high-resolution dataset captured with 128 channel Ouster LiDAR covering $4.2\,\text{km}^{2}$ of central Perth in Western Australia. The dataset comprises 144,000 LiDAR frames along with synchronized and georeferenced inertial and positioning data.
II Data Collection
@@
II-A Survey Objectives and Study Area Definition
@@
II-B Sensor Specifications
The sensing system comprises a 128-channel Ouster OS1 LiDAR and a RUG-3-IMX-5-Dual GNSS–IMU module [32]. The OS1 is a dual-return high-resolution scanner with a 45° vertical field of view, 0.35° angular resolution at 10 Hz, and a peak output of 2.6 million points per second. Each point includes 3D coordinates, intensity, and a time stamp, which supports detailed reconstruction of urban environments [33, 34]. The GNSS–IMU module integrates dual multi-frequency RTK receivers and a high-performance MEMS inertial unit, delivering sub-meter accuracy and full 6-DoF orientation. It supports GPS, GLONASS, Galileo, and BeiDou constellations. According to the manufacturer specifications, the GNSS operates at 5 Hz and the IMU at 73 Hz [35]. Unlike tightly coupled systems, the LiDAR and GNSS–IMU modules operate asynchronously, with synchronization and logging handled by a central acquisition system as detailed in subsection II-C.
II-C Sensor Setup and Hardware Configuration
@@
II-D Data Modalities, Formats, and Organization
@@
II-E Quality Assurance and Handling of Imperfections
@@ III Methods
III-A Multimodal Sensor Data Preprocessing
@@
III-A1 LiDAR Data
Raw LiDAR scans are decoded using the Ouster SDK [36], yielding point cloud frames with 3D coordinates, intensity, reflectivity, and ring ID attributes [33]. To enhance data quality, spurious returns from reflective or transparent surfaces are removed using statistical outlier filtering, which discards points deviating more than one standard deviation ($z=1.0$) from the mean distance within a $k$-nearest neighborhood (k=50k=50[37]. For efficiency and spatial consistency, voxel grid downsampling with a 0.05 m resolution is applied, replacing all points in each voxel with their centroid [38, 39]. Frames missing due to occlusion or transmission loss are reconstructed by linear interpolation between adjacent frames [40]. The processed point clouds are stored in .ply format with aligned timestamps for multimodal integration.
III-A2 GNSS Data
@@
$\hat{\mathbf{x}}_{t|t-1}=\hat{\mathbf{x}}_{t-1|t-1}+\int_{t-1}^{t}\mathbf{F}(\hat{\mathbf{x}}_{\tau},\mathbf{u}_{\tau})\,d\tau\vskip-2.84526pt$ (1)
@@ III-A3 IMU Data
III-B Multimodal Sensor Data Synchronization
@@
III-B1 LiDAR Velocity Estimation
To estimate LiDAR velocity, consecutive point cloud frames are registered using ICP algorithm, which yields a $\times4$ transformation matrix representing relative motion. The translational displacement $(\Delta x,\Delta y,\Delta z)$ extracted from this matrix is used to compute instantaneous speed as $v_{\text{LiDAR}}(t)=\frac{\sqrt{\Delta x^{2}+\Delta y^{2}+\Delta z^{2}}}{\Delta t}$, where $\Delta t$ is the time interval between frames [45].
III-B2 IMU Velocity Estimation
@@
$\vskip-5.69054ptv_{\text{IMU}}(t)=\left\|\mathbf{v}(t-\Delta t)+\int_{t-\Delta t}^{t}\mathbf{a}_{\text{corr}}(\tau)\,d\tau\right\|_{2}\vskip 5.69054pt$ (2)
@@
III-B3 Temporal Alignment with DTW
@@
III-B4 Temporal Alignment Methodology
@@
$\begin{split}D(i,j)&=\bigl(v_{\text{LiDAR}}(i)-v_{\text{IMU}}(j)\bigr)^{2}\\ &\quad+\min\{D(i-1,j),\,D(i,j-1),\,D(i-1,j-1)\}.\end{split}$ (3)
@@
III-C High-Precision 3D LiDAR Map Construction
@@
III-C1 Loop Registration and Pose Graph Optimization
@@
III-C2 GNSS-Based Loop Calibration
@@
$\mathbf{T}_{\text{loop}}^{\ast}=\arg\min_{\mathbf{T}}\sum_{i=1}^{M}\left\|\mathbf{G}_{\text{loop}}^{(i)}-\mathbf{T}\mathbf{T}_{\text{loop}}^{(i)}\right\|_{\mathbf{W}_{i}}^{2}\vskip 5.69054pt$ (5)
@@
III-C3 Global Merging Using ICP on Overlapping Regions
@@
III-D Three-Stage Assessment of 3D Map Alignment Accuracy
To assess the spatial accuracy of the proposed LIGMA framework, we implement a structured three-stage protocol. First, road centrelines are extracted from the reconstructed 3D LiDAR map and a georeferenced 2D base map from Google Maps, then compared to evaluate geometric alignment. Next, geometric alignment is assessed by computing point-wise correspondences between the LiDAR-derived and reference centrelines, with intersection regions excluded due to ambiguity in lane geometry and resolution. Lastly, intersection zones are analyzed separately using localized deviation metrics to capture alignment accuracy in these geometrically complex areas. These evaluation protocols are described in the following subsections.
III-D1 Extraction of Road Centrelines
@@
III-D2 Centreline-Based Alignment Error Metrics
@@
Figure 11: Qualitative comparison across six representative loops in Perth CBD (Seq 05, Seq 11, Seq 14, Seq 16, Seq 19, Seq 20). Top row: LiDAR-only reconstructions over satellite imagery (green). Middle row: reconstructions produced by the proposed LIGMA framework over the same areas (blue). Bottom row: centreline overlays showing the georeferenced reference (white), the LiDAR-only result (green), and LIGMA (red). The proposed method improves loop closure and intersection geometry while reducing drift and enhancing global consistency.
III-D3 Intersection-Based Local Geometric Consistency Metrics
@@
$\mathcal{E}_{\text{spatial}}=\frac{1}{K}\sum_{i=1}^{K}\left[\alpha\|p_{i}-m_{i}\|_{2}+\beta\cdot d_{\mathcal{T}}(p_{i},m_{i})\right]\vskip-5.69054pt$ (7)
@@
$\mathcal{E}_{\text{env}}=\frac{1}{K}\sum_{i=1}^{K}\max_{x\in\mathcal{B}(p_{i},r)}\min_{y\in\mathcal{B}(m_{i},r)}\|x-y\|_{2}\vskip-5.69054pt$ (8)
@@ IV-A Experimental Setup
Figure 13: Intersection centroid alignment using the LiDAR-only baseline. Green points represent the reference centroids and red points denote the LiDAR-only results. Yellow regions highlight areas of overlap between the two, emphasizing the degree of alignment.
Figure 14: Intersection centroid alignment using the proposed LIGMA framework. Green points represent the reference centroids and red points denote the fused results. Yellow regions indicate overlapping areas between the two, providing a clearer visualization of improved correspondence.
IV-B Effectiveness of Multimodal Sensor Data Synchronization
@@
IV-C High-Precision 3D LiDAR Map Construction Results
@@
IV-D Comparative Evaluation of Urban 3D Map Alignment
IV-D1 Visual Comparison of Mapping Results
TABLE I: Comparison of road centreline alignment metrics across all test sequences for the LiDAR-only baseline and the proposed framework. Relative improvements indicate percentage reductions in alignment error.
@@ IV-D1 Visual Comparison of Mapping Results
IV-D2 Loop-Based Centreline Comparison
To assess route-level consistency, we selected six representative loops from the twenty-one sequences in the Perth CBD (Seq 05, 11, 14, 16, 19, and 20). These loops span long straight segments, curved roads, and dense junctions. Figure 11 consolidates the visual evidence: for each loop, the top row shows the LiDAR-only baseline overlaid on satellite imagery, the middle row shows the reconstruction produced by LIGMA, and the bottom row presents the corresponding road centreline comparison against a 2D cartographic reference. The baseline typically exhibits drift-induced shear and irregular loop closures, with noticeable misalignment at intersections and along extended segments. In contrast, LIGMA yields coherent loop geometry and improved intersection alignment, with centrelines closely following the reference. While Fig. 11 illustrates typical behavior on six loops, the quantitative evaluation over all twenty-one sequences is summarized in Table I: our framework reduces centreline RMSE from $3.32$ m (LiDAR-only) to $1.24$ m on average, with consistent gains in the minimum, maximum, and standard deviation of the alignment error.
TABLE II: Intersection-level alignment metrics for 15 of 51 intersections (subset shown due to space constraints). CI: Centroidal Offset (m), MOD: Mean Orientation Deviation (°). Abs. = absolute reduction; Pct. = percentage reduction relative to baseline.